Solve equation, check answer. This is a radical equation with power functions.
x^(2/3) - x^(1/3) - 6 = 0

Answer by ewatrrr(24785)


Hi
solving
x^(2/3) - x^(1/3) - 6 = 0
Let a = x^(1/3)
a^2 - a - 6 = 0
factoring
(a-3)(a+2) = 0
(a-3)= 0 a = 3 x = 27 |a = x(1/3) OR a^3 = x
(a+2) = 0 a =-2 x = -8
CHECKING our Answers***
9 - 3 - 6 = 0 |x = 27
4 + 2 - 6 = 0 |x = -8